OpenStudy (anonymous):

what is the average rate of change from x=-10 to x=5?

OpenStudy (anonymous):

Well, you left out the function, but I'll leave the details to you. The average rate of change for a function \(f(x)\) over the interval \([a,b]\) is \[\frac{f(b)-f(a)}{b-a}\] where \(a=-10\) and \(b=5\).

OpenStudy (anonymous):

so would it be f(-5)/ 15 ?

OpenStudy (anonymous):

No. \[\text{ave. rate of change}=\frac{f(5)-f(-10)}{5-(-10)}=\frac{f(5)-f(-10)}{15}\]

OpenStudy (anonymous):

Subtraction *does not* work like you think it does with a general function \(f\). What is \(f(x)\)?

OpenStudy (anonymous):